# Cosmological simulations of the generation of cluster-scale radio emission from turbulent re-acceleration

### v1

@@ 2.2 Simulations of cosmic-ray electron spectra: CRATER & ROGER
Figure 3: Energy distributions of the tracers inside $R_{500}$ at two different redshifts, marking an epoch before ($10.86$ Gyr) and after ($12.65$ Gyr) the last importan merger in the system, and the onset of diffuse radio emission. Top-left: distribution of the ratio between the CR energy and the turbulent energy. The ratio increases with time and broadens after the merger; top-right: distribution of the CRe energy; bottom-left: kinetic, or turbulent, energy distribution; bottom-right: the gas energy distribution.
@@ 3.1 Cluster evolution
In the first phase of its evolution, up to $7\rm\leavevmode\nobreak\ Gyr$ since the start of the simulation, the cluster has accreted a large population of smaller clumps. Its shape is still asymmetric and perturbed. The merger between $\sim 8$ and $\sim 9\rm\leavevmode\nobreak\ Gyr$ finally produces a massive cluster in this region, with a mass that is approximately $\sim 85\%$ of the total final mass of the cluster at $z=0.0$.
From this point on, the cluster will still undergo several minor mergers, and a larger one, approximately with a mass ratio between 1:5 to 1:6. The latter is the most powerful event in terms of the injection of turbulence and for the production of diffuse radio emission. First a small satellite (which was entered the cluster about $1\rm\leavevmode\nobreak\ Gyr$ before) hits the cluster core at about $10.6\rm\leavevmode\nobreak\ Gyr$, and it creates asymmetries in the gas distribution, and it launches merger shock waves through the intracluster medium.
Again, between $\sim 11$ and $\sim 12\rm\leavevmode\nobreak\ Gyr$, the last major merger in the cluster launches several shocks. Soon after that, this merger distributes a large fraction of the gas from the inner part of the cluster to larger radii. At the same time this event generates significant turbulence. The formation of radio emission in our model, including the MegaRH-like structure out to $\rm R_{500}$, is closely related to this event.
@@
Figure 6: Radial profiles of the cluster at four different epochs along the Y line of sight, with the two different colors marking the Classical and MegaRH region.
@@ 3.2 Evolution of relativistic electrons
The population of relativistic electrons assigned to the Lagrangian tracers is evolved subject to radiative/collisional losses, adiabiatic terms, and additional (re)acceleration by shocks and turbulence. We present results for three models (”C”,”CT” and ”CST”) already introduced in Sec. 2.2. Model ”CST” is the baseline model, which contains all physical losses and plausible (re)acceleration scenarios expected in the ICM. Here, all CRe are injected at $z=2$, with an initial fraction $\phi=10^{-6}$ with respect the density of thermal electrons, and using $\eta_{B}=0.035$ to estimate the unresolved small-scale dynamo onboard of tracers.
@@
The evolving momentum spectra of all CRe in the simulated volume are shown in Fig. 2, where we consider the same snapshots of Fig. 1, with the addition of the initial momentum spectra of all tracers, one time step after the injection of the input power-law distribution. Early on, when the cluster mass is still composed of several approaching substructures, the CRe spectra mainly suffer Inverse Compton losses. At $z\approx 1$ the difference between the three models is not very large. Still, the continuous effect of turbulent re-acceleration has produced a significant excess of high momentum electrons ($\gamma\geq 10^{3}$). When the additional contribution from merger shocks is included, a small fraction of the CRe population reaches $\gamma\geq 10^{5}$. The actual number of CRe injected by shocks crucially depends here on the assumed acceleration efficiency by DSA, which is highly uncertain (e.g. Bykov et al., 2019; Böss et al., 2023b). However, for all plausible choices of the acceleration efficiency as a function of Mach number, the bulk of the high-energy population of CRe depends on turbulent re-acceleration acting on the initial seed population and not on the freshly accelerated particles by DSA. This is shown by the small difference between the CT and the CST models in most snapshots. In both cases, the plots show a steady growth of the momentum distribution , namely a large fraction of the momentum distribution has $\gamma\geq 10^{3}$, through the approximate balance between turbulent re-acceleration, compression and losses, as studied already in Paper I.
@@
Starting from approximately $z\sim 1$, recurring mergers sustain large levels of turbulence within $\rm R_{500}$, injecting a budget of total turbulent kinetic energy up to $\sim 2\cdot 10^{61}\rm\ erg$. This turbulent budget, in turn, leads to an increase in CRe energy (both in the CT and CST models), which is key for the formation of extended radio sources. Figure 3 shows the ratio between the CRe energy and the kinetic energy measured for all tracers in the MegaRH region, shortly before and after the last important merger experienced by the cluster. The distribution of CRe energy gets significantly wider after the merger, following the re-acceleration by the turbulence injected in this region. In the $\sim 1.8\rm\leavevmode\nobreak\ Gyr$ time difference considered here, the differences in the CRe distribution, or in the ratio between CRe energy and kinetic energy remain significantly larger than the differences of the distribution of turbulent kinetic energy across the same time difference. This is consistent with the fact (already suggested in our Paper I) that a large fraction of CRe in this regime can reach a long-term balance between gain and losses, while in the same time a significant fraction of the turbulent kinetic energy gets thermalised (consistent with the increase of thermal energy measured over the same timescale.
@@
Figure 7: Top: Radio maps of the cluster at 10.61 and 12.76 Gyr, along three lines of sight. Bottom: radial profiles of the surface brightness for the three lines of sight. The $z$ line-of-sight roughly corresponds to the merger axis of the latest major merger in the cluster.
@@ 3.3 Evolution of radio emission
We also compute the radial profile of the surface brightness, by radially averaging the emission within fixed radial annuli from the cluster centre, following a procedure similar to the MegaRH discovery paper by Cuciti et al. (2022). The cluster centre is selected as the brightest pixel of the projected image, as is done in observations. The radial profile of ClassicalRH is usually taken to follow an exponential profile (e.g. Murgia et al., 2009): $\rm I(r)=I_{0}e^{-r/r_{e}}$, where $\rm r_{e}$ is the $e$-folding radius and $\rm I_{0}$ is the maximum of the luminosity. The fit is executed as follows: First, we performed a preliminary interpolation of the radial surface brightness profile of radio emission with the exponential function. Next, the best fit parameters are used to create a new point symmetric image, by means of the radial profile function. This image is then convolved with a Gaussian beam which $\sigma=FWHM/2.35$ of LOFAR HBA. Finally, this image is radially averaged and the best-fit parameters are derived using a least-square method. Cuciti et al. (2022) calculated radial profiles by integrating over concentric annuli (including emission below $3\sigma$), and computing the noise as $\rm rms/\sqrt{\rm N_{beams}}$, where $\rm N_{beams}$ is the number of beams within each annulus. The rms is noise calculated from the images where the MegaRH have been detected. Considering that our simulated MegaRH would be too dim for detection, we refrain from performing such an analysis here. Our main goal is to assess whether a second, extended component in the surface brightness emerges during the merger, albeit at a low emission level considering the low mass of this system. We also compute the average spectral index between 50 MHz and 150 MHz within the same concentric annuli.
@@
Figure 5 shows a longer sequence of evolutionary steps as well as the radial profiles of magnetic fields and turbulent velocities. Also shown is the radial profile of the spectral index of radio emission between 50 and 150 MHz. Already soon after the first major merger leading to the cluster’s final mass ($t\sim 8.58\rm\leavevmode\nobreak\ Gyr$ in Fig. 5) the cluster forms a centrally located, smooth radio profile, compatible with a low-power and ultra-steep spectrum radio halo. This source remains detectable for less then 1 Gyr, due to the sudden increase of turbulence following the merger of the parent satellites. This is followed by a phase of low, or absent, radio luminosity for the subsequent 1.5 Gyr, caused by a drop in turbulence and marked by a very steep radio spectral index ($\alpha\leq-1.75$).
@@
After the first merger, the entrance of a small satellite again boosts the diffuse radio emission above the detection limit. Its effect is observed in the radial profile as a sudden increase of flux in the sector swept by the satellite ($t\sim 10.61\rm\leavevmode\nobreak\ Gyr$ in Fig. 5)). This event creates a deformation of the profile which does not resemble a MegaRH. However, it still produces significant re-acceleration that enables the formation of extended radio emission at a later stage. Moreover, the launching of merger shock waves at this epoch promotes the formation of a radio relic-like emission feature, powered by shock (re)acceleration in the northern sector of the cluster. This is also visible in the emergence of a straight power-law in the momentum spectra of Fig. 2 for the same epoch ($z=0.27$), in our CST model which also includes the shock (re)acceleration of electrons via DSA as explained in Sec. 2, (e.g. Inchingolo et al., 2022; Smolinski et al., 2023).
The shape of the profile follows the collision geometry (see Fig. 6), as can be seen from the asymmetric emission patterns visible in the radio map of Fig. 4.
@@
After another $100\rm\leavevmode\nobreak\ Myr$ (i.e. around 12.7 Gyr) we can clearly observe the formation of a second component of the radial profile of surface brightness, which follows from the asymmetric distribution of radio emission boosted by the turbulence in the wake of the accreted subcluster.
@@
Our simulations clearly show that turbulence injected by the last major merger is important for creating, both, a ClassicalRH and a MegaRH-like source. Two closely related effects contribute to this: a) the merger generates a significant turbulent wake along the merger axis, with rms velocities comparable to those injected in the cluster centre, and b) the large-scale motions induced by the merger displace a large pool of relativistic electrons that are already re-energised by past acceleration cycles. After 800 Myr, such re-energised electrons get dispersed to more peripheral cluster regions, i.e. as far as $R_{\rm 200}$, while being still barely detectable in radio. In both cases, these mechanisms tend to produce a rather asymmetric emission towards the external cluster regions, which is at variance with about a half of known MegaRH (Cuciti et al., 2022). However, the actual luminosity changes with time as well as with the direction of the observation.
@@
3.4 How do Mega Radio Halos form?
@@
In the simulation, the fundamental mechanism responsible for the formation of diffuse radio emission, both, in the form of Classical and Mega RHs, is CR re-acceleration by turbulence. This turbulence is injected on scales of $\sim\rm Mpc$ during mergers. This turbulence is more solenoidal and more subsonic in the ClassicalRH regions, than in the MegaRH, where it is more compressive and has a higher Mach number.
During the post-merger phase, we observe a significant re-energisation of electrons. This results in the appearance of a second component in the radial brightness profiles, as is observed in real MegaRH emission. Re-energisation is primarily triggered by the significant increase in turbulent velocity and a comparatively slower rise in the magnetic field strength. The radio emitting electrons of the MegaRH were located in the cluster centre before the last merger, and got advected to outer radii (and encountered significantly different turbulent conditions) only their last $\sim\rm\leavevmode\nobreak\ Gyr$ of evolution (Beduzzi et al., 2023). This process allows a large fraction of CRe to keep a $\gamma\gtrsim 10^{3}$ energy level, which makes them subject to stochastic turbulent re-acceleration for timescales longer than their cooling time (consistent with what anticipated in Paper I).
@@
We stress that under realistic observing conditions, our simulated MegaRH would hardly be detectable. Its low mass makes the whole process subcritical to produce radio emission beyond the LOFAR HBA and LBA detection limit, if we place the system at $z\approx 0.1$. In principle, the results of this study can be qualitatively extrapolated to the regime of higher-mass clusters, considering the approximate self-similar dependence between the turbulent kinetic energy and the cluster mass, i.e. $E_{\rm turb}\propto M^{5/3}$ (e.g. Cassano & Brunetti, 2005; Vazza et al., 2006), and the approximate scaling between the radio luminosity and the kinetic power, i.e. $L_{R}\propto E_{\rm turb}/t_{\rm eddy}$, where $t_{\rm eddy}\sim L/\sigma_{v}$ is the eddy turnover time for an injection scale L (in turn, typically a fraction of the cluster virial radius, e.g. Brunetti & Vazza, 2020). However, other processes are likely not to obey similar self-similar rescaling relations, e.g. the injection scale of turbulence, the damping of MHD waves as well as the ratio between synchrotron losses and Inverse Compton losses. For this reason, it is presently uncertain how to extrapolate these results to clusters with 10 times larger mass, which are closer to the real observations. This will be subject of forthcoming work.
Figure 9: Radial profiles of radio surface brightness at 150MHz, considering four different injection times for the seed CRe, and for two different epochs of observations ($z=0.151$ and $z=0.027$).

### v2

@@ 2.2 Simulations of cosmic-ray electron spectra: CRATER & ROGER
Figure 3: Energy distributions of the tracers inside $R_{500}$ at two different redshifts, marking an epoch before ($10.86$ Gyr) and after ($12.65$ Gyr) the last important merger in the system, and the onset of diffuse radio emission. Top-left: distribution of the ratio between the CR energy and the turbulent energy. The ratio increases with time and broadens after the merger; top-right: distribution of the CRe energy; bottom-left: kinetic, or turbulent, energy distribution; bottom-right: the gas energy distribution.
@@ 3.1 Cluster evolution
In the first phase of its evolution, up to $7\rm\leavevmode\nobreak\ Gyr$ (panel 1a in the figure) since the start of the simulation, the cluster has accreted a large population of smaller clumps. Its shape is still asymmetric and perturbed. The merger between $\sim 8$ and $\sim 9\rm\leavevmode\nobreak\ Gyr$ (panels 1b to 1d) finally produces a massive cluster in this region, with a mass that is approximately $\sim 85\%$ of the total final mass of the cluster at $z=0.0$.
From this point on, the cluster will still undergo several minor mergers, and a larger one, approximately with a mass ratio between 1:5 to 1:6 (we analyse the role of the merger mass ratio in Sec.4). The latter is the most powerful event in terms of the injection of turbulence and for the production of diffuse radio emission. First a small satellite (which was entered the cluster about $1\rm\leavevmode\nobreak\ Gyr$ before) hits the cluster core at about $10.6\rm\leavevmode\nobreak\ Gyr$, and it creates asymmetries in the gas distribution, and it launches merger shock waves through the ICM. Again, between $\sim 11$ and $\sim 12\rm\leavevmode\nobreak\ Gyr$, the last major merger in the cluster launches several shocks (panel 1f). Soon after that, this merger distributes a large fraction of the gas from the inner part of the cluster to larger radii. At the same time this event generates significant turbulence. The formation of radio emission in our model, including the MegaRH-like structure out to $\rm R_{500}$, is causally connected to this event, but the radio emission peaks a few hundreds of million years later than this, owing to the fact that turbulent energy cascade only after an eddy turnover time. Moreover, the most spectacular effect here is in the production of extended diffuse emission beyond the innermost $\sim\rm Mpc^{3}$ occupied by the Classical RH emission, which we shall discuss in the next Sections.
@@
Figure 6: Radial profiles of the cluster at four different epochs along the Y line of sight. The region used for the fit where selected as in Cuciti et al. (2022) and in the figures it is find between the centre and the vertical dashed line.
@@ 3.2 Evolution of relativistic electrons
The population of relativistic electrons assigned to the Lagrangian tracers is evolved subject to radiative and collisional losses, adiabiatic terms, and additional (re-)acceleration by shocks and turbulence. We present results for three models (”C”,”CT” and ”CST”) already introduced in Sec. 2.2. Model ”CST” is the baseline model, which contains all physical losses and plausible (re-)acceleration scenarios expected in the ICM. Here, all CRe are injected at 3.37 Gyr ($z\approx 2$), with an initial fraction $\phi=10^{-6}$ with respect the density of thermal electrons, and using $\eta_{B}=0.035$ to estimate the unresolved small-scale dynamo on board of tracers.
@@
The evolving momentum spectra of all CRe in the simulated volume are shown in Fig. 2, where we consider the same snapshots of Fig. 1, with the addition of the initial momentum spectra of all tracers, one time step after the injection of the input power-law distribution. Early on, when the cluster mass is still composed of several approaching substructures, the CRe spectra mainly suffer Inverse Compton losses. At 5.96 Gyr ($z\approx 1$) the difference between the three models is not very large. Still, the continuous effect of turbulent re-acceleration has produced a significant excess of high momentum electrons ($\gamma\geq 10^{3}$). When the additional contribution from merger shocks is included, a small fraction of the CRe population reaches $\gamma\geq 10^{5}$. The actual number of CRe injected by shocks crucially depends here on the assumed acceleration efficiency by DSA, which is highly uncertain (e.g. Bykov et al., 2019; Böss et al., 2023b). However, for all plausible choices of the acceleration efficiency as a function of Mach number, the bulk of the high-energy population of CRe depends on turbulent re-acceleration acting on the initial seed population and not on the freshly accelerated particles by DSA. This is shown by the small difference between the CT and the CST models in most snapshots. In both cases, the plots show a steady growth of the momentum distribution , namely a large fraction of the momentum distribution has $\gamma\geq 10^{3}$, through the approximate balance between turbulent re-acceleration, compression and losses, as studied already in Paper I.
@@
Starting from approximately $8\leavevmode\nobreak\ \rm Gyr$, recurrent mergers sustain large levels of turbulence within $\rm R_{500}$, and inject a significant budget of turbulent kinetic energy. By summing all the recorded turbulent kinetic energy by our tracers, we measure up to $2\cdot 10^{62}$ erg of solenoidal turbulent kinetic energy injected in the cluster during the assembly of most of the cluster mass, followed by an overall injection of $3\cdot 10^{61}$ erg at the end of the simulation. This turbulent budget, in turn, leads to an increase in CRe energy (both in the CT and CST models), which is key for the formation of extended radio sources. Figure 3 shows the ratio between the CRe energy and the kinetic energy measured for all tracers in the MegaRH region, shortly before and after the last important merger experienced by the cluster. The distribution of CRe energy gets significantly wider after the merger, following the re-acceleration by the turbulence injected in this region. In the $\sim 1.8\rm\leavevmode\nobreak\ Gyr$ time difference considered here, the differences in the CRe distribution, or in the ratio between CRe energy and kinetic energy remain significantly larger than the differences of the distribution of turbulent kinetic energy across the same time difference. This is consistent with the fact (already suggested in our Paper I) that a large fraction of CRe in this regime can reach a long-term balance between gain and losses, while in the same time a significant fraction of the turbulent kinetic energy gets thermalised (consistent with the increase of thermal energy measured over the same timescale). Based on the quantities recorded by tracers, we measure that the solenoidal turbulent kinetic energy is initially $\sim 19\%$ of the total energy after the assembly of the bulk of the cluster mass, and it decreases to $\sim 4\%$ by the end of the simulation, marking a gradual thermalisation of the kinetic energy induced by accretion events. During the same time, however, we measure a steady increase in the estimated total energy of all CRe in our tracers, which amount to $6\cdot 10^{-5}\%$ of the total gas energy at $\sim 8\leavevmode\nobreak\ \rm Gyr$ and finish having $4\cdot 10^{-4}\%$ of the total gas energy towards the end of the simulation.
@@
Figure 7: Representations of both the projection of the cluster and its radial profile along that direction at the same ages. Top: Radio maps of the cluster at 10.61 and 12.76 Gyr, along three lines of sight. Bottom: radial profiles of the surface brightness for the three lines of sight. The $z$ line-of-sight roughly corresponds to the merger axis of the latest major merger in the cluster. The region used to compute the best-fit for the radial profile was manually selected as in Cuciti et al. (2022), and is delimited by the vertical dashed line in the plot.
@@ 3.3 Evolution of radio emission
We also compute the radial profile of the surface brightness, by radially averaging the emission within fixed radial annuli from the cluster centre, following a procedure similar to the MegaRH discovery paper by Cuciti et al. (2022). The cluster centre is selected as the brightest pixel of the projected image, as is done in observations. The radial profile of ClassicalRH is usually taken to follow an exponential profile (e.g. Murgia et al., 2009): $\rm I(r)=I_{0}e^{-r/r_{e}}$, where $\rm r_{e}$ is the $e$-folding radius and $\rm I_{0}$ is the maximum of the luminosity. The fit is executed as follows. First, we performed a preliminary interpolation of the radial surface brightness profile of radio emission with the exponential function, within a manually selected radial range from the peak of radio emission. Although this procedure is semi-empiric and it does not rely on an automatic selection of the radii to associate with a ClassicalRH or a MegaRH region. This is consisent with the procedure followed in the discovery paper of MegaRH, whose logic we want to reproduce here. A future larger samples of simulated and observed MegaRH will infer a proper statistics on the extension of this sources, enabling us to design automated procedures for the identification of this structures, but this goal goes beyond the purpose of this paper. Next, the best fit parameters are used to create a new point symmetric image, by means of the radial profile function. This image is then convolved with a Gaussian beam which $\sigma=FWHM/2.35$ of LOFAR HBA. Finally, this image is radially averaged and the best-fit parameters are derived using a least-square method. Cuciti et al. (2022) calculated radial profiles by integrating over concentric annuli (including emission below $3\sigma$), and computing the noise as $\rm rms/\sqrt{\rm N_{beams}}$, where $\rm N_{beams}$ is the number of beams within each annulus. The rms is noise calculated from the images where the MegaRH have been detected. Considering that our simulated MegaRH would be too dim for detection, we refrain from performing such an analysis here. Our main goal is to assess whether a second, extended component in the surface brightness emerges during the merger, albeit at a low emission level considering the low mass of this system. We also compute the average spectral index between 50 MHz and 150 MHz within the same concentric annuli.
@@
Figure 5 shows a longer sequence of evolutionary steps as well as the radial profiles of magnetic fields and turbulent velocities. Also shown is the radial profile of the spectral index of radio emission between 50 and 150 MHz. Already soon after the first major merger leading to the cluster’s final mass ($t\sim 8.58\rm\leavevmode\nobreak\ Gyr$ in Fig. 5, 5a) the cluster forms a centrally located, smooth radio profile, compatible with a low-power and ultra-steep spectrum radio halo. This source remains detectable for less then 1 Gyr, due to the sudden increase of turbulence following the merger of the parent satellites. This is followed by a phase of low, or absent, radio luminosity for the subsequent 1.5 Gyr, caused by a drop in turbulence and marked by a very steep radio spectral index ($\alpha\leq-1.75$).
@@
After the first merger, the entrance of a small satellite again boosts the diffuse radio emission above the detection limit. Its effect is observed in the radial profile as a sudden increase of flux in the sector swept by the satellite ($t\sim 10.61\rm\leavevmode\nobreak\ Gyr$ in Fig. 5, 5c). This event creates a deformation of the profile which does not resemble a MegaRH. However, it still produces significant re-acceleration that enables the formation of extended radio emission at a later stage. Moreover, the launching of merger shock waves at this epoch promotes the formation of a radio relic-like emission feature, powered by shock (re-)acceleration in the northern sector of the cluster. This is also visible in the emergence of a straight power-law in the momentum spectra of Fig. 2 at 10.7 Gyr, in our CST model which also includes the shock (re-)acceleration of electrons via DSA as explained in Sec. 2, (e.g. Inchingolo et al., 2022; Smolinski et al., 2023). The shape of the profile follows the collision geometry (see Fig. 6), as can be seen from the asymmetric emission patterns visible in the radio map of Fig. 4.
@@
After another $100\rm\leavevmode\nobreak\ Myr$ (i.e. around 12.65 Gyr in Fig. 6) we can clearly observe the formation of a second component of the radial profile of surface brightness, which follows from the asymmetric distribution of radio emission boosted by the turbulence in the wake of the accreted subcluster.
@@
Our simulations clearly show that turbulence injected by the last major merger is important for creating, both, a ClassicalRH and a MegaRH-like source. Two closely related effects contribute to this: the merger generates a significant turbulent wake along the merger axis, with rms velocities comparable to those injected in the cluster centre, and the large-scale motions induced by the merger displace a large pool of relativistic electrons that are already re-energised by past acceleration cycles. After 800 Myr, such re-energised electrons get dispersed to more peripheral cluster regions, i.e. as far as $R_{\rm 200}$, while being still barely detectable in radio. In both cases, these mechanisms tend to produce a rather asymmetric emission towards the external cluster regions, which is at variance with about a half of known MegaRH (Cuciti et al., 2022). However, the actual luminosity changes with time as well as with the direction of the observation.
@@
3.4 Exploring the formation of mega radio halos
@@
In the simulation, the fundamental mechanism responsible for the formation of diffuse radio emission, both, in the form of Classical and Mega RHs, is CR re-acceleration by turbulence. This turbulence is injected on scales of $\sim\rm Mpc$ during mergers. This turbulence is more solenoidal and more subsonic in the ClassicalRH regions, than in the MegaRH, where it is more compressive and has a higher Mach number. Indeed based on the gas temperature profile in this region at $\sim 12.65\leavevmode\nobreak\ \rm Gyr$(see also Paper I), the local sound speed is $c_{s}\sim 960\rm\leavevmode\nobreak\ km/s$ on average in the Classical RH region, while it is $c_{s}\sim 750\rm\leavevmode\nobreak\ km/s$ in the MegaRH region; at the same time, the rms velocity related to the solenoidal component is $\sigma_{v}\sim 280\leavevmode\nobreak\ \rm km/s$ in the ClassicalRH region and $\sigma_{v}\sim 250\leavevmode\nobreak\ \rm km/s$ in the MegaRH region, corresponding to a turbulent Mach number $\mathcal{M}_{t}\sim 0.29$ in the first case, and to $\mathcal{M}_{t}\sim 0.33$ in the second case. During the post-merger phase, we observe a significant re-energisation of electrons. This results in the appearance of a second component in the radial brightness profiles, as is observed in real MegaRH emission. Re-energisation is primarily triggered by the significant increase in turbulent velocity and a comparatively slower rise in the magnetic field strength. The radio emitting electrons of the MegaRH were located in the cluster centre before the last merger, and got advected to outer radii (and encountered significantly different turbulent conditions) only their last $\sim\rm\leavevmode\nobreak\ Gyr$ of evolution (Beduzzi et al., 2023). This process allows a large fraction of CRe to keep a $\gamma\gtrsim 10^{3}$ energy level, which makes them subject to stochastic turbulent re-acceleration for timescales longer than their cooling time (consistent with what anticipated in Paper I).
@@
We stress that under realistic observing conditions, our simulated MegaRH would hardly be detectable. Its low mass makes the whole process subcritical to produce radio emission beyond the LOFAR HBA and LBA detection limit, if we place the system at $z\approx 0.1$. In principle, the results of this study can be qualitatively extrapolated to the regime of higher-mass clusters, considering the approximate self-similar dependence between the turbulent kinetic energy and the cluster mass, i.e. $E_{\rm turb}\propto M^{5/3}$ (e.g. Cassano & Brunetti, 2005; Vazza et al., 2006), and the approximate scaling between the radio luminosity and the kinetic power, i.e. $L_{R}\propto E_{\rm turb}/t_{\rm eddy}$, where $t_{\rm eddy}\sim L/\sigma_{v}$ is the eddy turnover time for an injection scale L (in turn, typically a fraction of the cluster virial radius, e.g. Brunetti & Vazza, 2020) and $\sigma_{v}$ is the rms velocity dispersion computed within the same scale. However, other processes are likely not to obey similar self-similar rescaling relations, e.g. the injection scale of turbulence, the damping of MHD waves as well as the ratio between synchrotron losses and Inverse Compton losses. For this reason, it is presently uncertain how to extrapolate these results to clusters with 10 times larger mass, which are closer to the real observations. This will be subject of forthcoming work.
Figure 9: Radial profiles of radio surface brightness at 150MHz, considering four different injection times for the seed CRe. The injection epochs were chosen to bracket the important moments marking the assembly of most of the cluster mass, at $t\approx 8.6\rm\leavevmode\nobreak\ Gyr$, as well as the last major merger experienced by the cluster, at $t\approx 12\rm Gyr$ (see also Fig.1).